A 0.40 mm diameter copper wire carries a current of

3μA. Find the current density and he electron drift density in the wire


1
Expert's answer
2022-01-31T10:43:13-0500

Now current density

"J=\frac{I}{A}\\J=\frac{I}{\pi r^2}"

"r=\frac{D}{2}=0.2mm"


"A=\pi r^2=3.14\times(0.2\times10^{-3})^2=3.14\times4\times10^{-8}"

"J=\frac{3\times10^{-6}}{3.14\times4\times10^{-8}}=23.88A/m^2"

Drift velocity

"V_d=\frac{J}{ne}"


"V_d=\frac{23.88}{8.49\times10^{28}\times1.6\times10^{-19}}=4.5\times10^{-9}m/sec"